Suppose demand for wind turbines is Q = 110-3P, where P is the price. The dominant producer in this industry is “Winnie’s Wind Turbines”. There are also a number of small price-taking firms that can be represented by the supply function S(P)=P-10. The marginal cost of production for the dominant firm is given by mcd=10 and the total cost function is given by 10qd.

What quantity would Winnie’s Wind Turbines supply in the wind turbine market?

What would be the market price for a wind turbine?

What quantity would the fringe supply?
Answer #1

Market demand is Q = 110 – 3P

Fringe supply is Q = P – 10

Residual demand = 110 – 3P – P + 10

= 120 – 4P

Inverse residual demand is P = 30 – 0.25Q

Marginal revenue is MR = 30 – 0.5Q

Use MR = MC

30 – 0.5Q = 10

Q = 20/0.5 = 40 units

P = $20

Fringe firms will supply 20 – 10 = 10 units

Thus,

Winnie’s Wind Turbines would supply 40 units in the wind turbine market

The price would be $20 for a wind turbine

Fringe supply 10 units